[Hallc_running] [New Logentry] RC daily report

berdnik at jlab.org berdnik at jlab.org
Thu Nov 18 08:55:01 EST 2021

Logentry Text:

Day shift yesterday (Nov 17) almost no beam due separator issues. 
Swing shift beam delivery started at 18-00, quality was bad, lots of FSD trips, current less then required 70uA. 
MCC took beam away again around 20-30 for 45 minutes to steer A2 apperture, as result stable 70uA beam operation overnight and morning.     
At 8-45 am we have 6642 event out of 20k. The plan is continue with data taking when available. MCC will take beam away around 10-30 top continue work on separator, estimated down time 2hours, may be more depending on situation. 

Run plan:

Data taking LH2 till Friday(Nov 19) day shit, then rotate SHMS to 6.6 degree and continue with data taking over the weekend.

Pending issues:

HMS dipole NMR probe dying https://logbooks.jlab.org/entry/3945416 and https://logbooks.jlab.org/entry/3945477 

RC meeting:

Let's have a meeting today at 10:00 AM EST.
Connection information: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3945821
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211118/d52243c2/attachment.html>

More information about the Hallc_running mailing list